Circular No. 7106

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999Y IN ANONYMOUS GALAXY D. Reiss, Department of Astronomy, University of Washington, re-

ports the discovery by the Mount Stromlo Abell Cluster Supernova Search Team (cf. IAUC 6639) of an apparent supernova (V = 19.0) on V and R CCD images (limiting mag V = 21.0) taken on Jan. 10 by S. Sabine at the Mount Stromlo 1.27-m telescope (+ Macho Camera). SN 1999Y is located near Abell 3142 at α = 3^{h}36^{m}55^{s}.81, δ = −39^{o}33′04′′.9 (equinox 2000.0), which is 3′′.8 west and 0′′.9 south of the host galaxy’s center. A nearby star has position end figures 57^{s}.39, 51′′.7. SN 1999Y was not visible on similar images taken on Dec. 23 (limiting mag V = 21.0), but it was confirmed by D. Reiss and L. Germany with the Cerro Tololo 0.9-m reflector in 120-s exposures (in B, V, R, and I) obtained on Jan. 13 and 14, in which it appeared ∼ 0.3 mag brighter. Finding charts are available

at http://msowww.anu.edu.au/~reiss/Abell SNSearch.

SUPERNOVA 1999Z IN UGC 5608 M. Schwartz, Cottage Grove, OR, reports his discovery of an apparent

supernova (mag 16.2) on an unfiltered CCD frame taken on Feb. 8.437 UT with the Tenagra I 0.35-m internet-controlled supernova patrol telescope in Sonoita, AZ. SN 1999Z is located at α = 10^{h}22^{m}37^{s}.23, δ = +27^{o}21′19′′.8 (equinox 2000.0), which is ∼ 3′′.6 west and 9′′ south of the nucleus of UGC 5608. The new object was confirmed on a frame taken on Feb. 9.153. The object does not appear on either the first or the second Palomar Sky Survey scans, and it is absent from the USNO A1.0 catalogue. P. Zhao confirms the existence of this star on Feb. 9.366 from an observation made with the F. L. Whipple Observatory 1.2-m telescope (+ AndyCam + BVRI filters).

SUPERNOVAE 1998es AND 1999D E. Prosperi, Larciano, Italy, reports the following CCD magnitudes

(cf. IAUC 7094): SN 1998es in NGC 632, Feb. 6.747 UT, R = 15.2; SN 1999D in NGC 3690, Feb. 6.998, R = 15.8.

COMET C/1998 M5 (LINEAR) Visual m_{1} estimates: Jan. 4.98 UT, 9.5 (J. Bortle, Stormville, NY,

0.41-cm reflector); 20.78, 9.3 (M. Meyer, Frauenstein, Germany, 20×100 binoculars); Feb. 5.78, 9.7 (R. J. Bouma, Groningen, The Netherlands, 0.25m refl.); 9.18, 9.7 (K. Hornoch, Lelekovice, Czech Rep., 10×80 binoculars).

1999 February 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green